# GCC: A 3DGS Inference Architecture with Gaussian-Wise and Cross-Stage Conditional Processing

### v1

@@ Abstract.
3D Gaussian Splatting (3DGS) has emerged as a leading neural rendering technique for high-fidelity view synthesis, prompting the development of dedicated 3DGS accelerators for mobile applications. Through in-depth analysis, we identify two major limitations in the conventional decoupled preprocessing-rendering dataflow adopted by existing accelerators: 1) a significant portion of preprocessed Gaussians are not used in rendering, and 2) the same Gaussian gets repeatedly loaded across different tile renderings, resulting in substantial computational and data movement overhead. To address these issues, we propose GCC, a novel accelerator designed for fast and energy-efficient 3DGS inference. At the dataflow level, GCC introduces: 1) cross-stage conditional processing, which interleaves preprocessing and rendering to dynamically skip unnecessary Gaussian preprocessing; and 2) Gaussian-wise rendering, ensuring that all rendering operations for a given Gaussian are completed before moving to the next, thereby eliminating duplicated Gaussian loading. We also propose an alpha-based boundary identification method to derive compact and accurate Gaussian regions, thereby reducing rendering costs. We implement our GCC accelerator in 28nm technology. Extensive experiments demonstrate that GCC significantly outperforms the state-of-the-art 3DGS inference accelerator, GSCore, in both performance and energy efficiency.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 58th IEEE/ACM International Symposium on Microarchitecture; October 18–22, 2025; Seoul, Koreaisbn: 978-X-XXXX-XXXX-X/XX/XX
@@ 1. Introduction
First, the two-stage “preprocess-then-render” approach introduces redundant Gaussian movements and computations. During the rendering phase, the early termination strategy in $\alpha$-blending results in the actual number of Gaussians used for rendering being significantly smaller than those within the view frustum. This implies that the 3D Gaussians corresponding to those non-rendered 2D ones can be entirely excluded during preprocessing, thereby reducing both data movements and computations. However, in the standard two-stage dataflow, all 3D Gaussians undergo preprocessing regardless of their actual utilization in subsequent rendering, introducing substantial redundancy. For example, our profiling reveals that approximately 40% of runtime in the state-of-the-art GSCore (Lee et al., 2024) accelerator is dedicated to preprocessing 3D Gaussians, while over 60% of the preprocessed Gaussians remain unused during subsequent rendering.
Second, the tile-wise rendering incurs repeated data accesses for the same Gaussians. At the beginning of rendering, all preprocessed 2D Gaussians establish indexing relationships with image tiles based on their spatial coordinates and bounding sizes. The rendering then processes tiles in scanline order: for each tile, it 1) retrieves all associated Gaussians, 2) sorts them by depth, and 3) computes $\alpha$-blending to determine RGB values for every pixel within the tile. For tiles of fixed size, when a Gaussian overlaps with multiple tiles, it means it will be repeatedly loaded during the rendering process of different tiles. This results in the actual number of Gaussian loading being significantly higher than the number of effective Gaussians in the entire rendering process, which is inefficient in terms of memory access.
@@ 3. GCC Dataflow
(7) $(\bm{p}-\bm{\mu^{\prime}})^{\top}\bm{\Sigma^{\prime}}^{-1}(\bm{p}-\bm{\mu^{\prime}})\leq 2\cdot\ln(255\cdot\omega)$
@@ Methods
1:Input: Pixel grid $P$, projected Gaussian center $\mu^{\prime}$, effective condition $E(p)$, visited map $S$ (initialized to 0)
@@
3:Let $p_{c}\leftarrow$ the nearest pixel to $\mu^{\prime}$ within image bounds.
4:Mark $S[p_{c}]\leftarrow 1$, set $P_{\text{influence}}\leftarrow\{p_{c}\}$, and initialize queue $Q\leftarrow\{p_{c}\}$.
5:while $Q\neq\emptyset$ do
6:     $p\leftarrow$ dequeue($Q$)
7:     for each neighbor $q$ of $p$ such that $q$ is within bounds and $S[q]=0$ do
8:         if $E(q)$ holds then
9:              Mark $S[q]\leftarrow 1$, add $q$ to $P_{\text{influence}}$, and enqueue $q$ into $Q$.
10:         else
11:              Mark $S[q]\leftarrow 1$triangleright Exclude pixels along the direction towards $q$
12:         end if
13:     end for
14:end while
15:return $P_{\text{influence}}$
@@ 4.4. Alpha Unit
In the 3DGS rendering pipeline, each projected Gaussian must compute a transparency weight $\alpha$ for its coverage on the image plane. As shown in Equation  7, $\alpha$ is modeled as the exponential of a negative quadratic form, derived from the pixel’s relative position and the inverse covariance matrix. To efficiently support this computation in hardware, we implement a lookup-table-based EXP unit using piecewise linear approximation. Given the range of meaningful $\alpha$ values lies in $[1/255,1)$, the corresponding exponent input is constrained to $[-5.54,0)$. Our LUT covers only this interval; inputs below $-5.54$ are clamped to $\alpha=0$, and those above 00 are saturated to $\alpha=1$. The LUT contains 16 linear segments, ensuring less than 1% approximation error while significantly reducing area and storage overhead. Unlike GSCore (Lee et al., 2024), which suffers from FP16 overflows in EXP operations due to value mapping, our design avoids such issues through fully fixed-point arithmetic in the EXP unit.
@@ 4.5. Blending Unit
Since 3DGS requires strict back-to-front blending, the pipeline enforces ordering at the block level. If a later Gaussian in the sorted sequence attempts to access a block whose previous Gaussian has not yet completed processing, the pipeline stalls until the prior Gaussian finishes, ensuring correctness of blending results. To avoid unnecessary computation, we introduce a transmittance mask ($T_{\text{mask}}$) that disables blending for blocks whose $T$ value has already fallen below a predefined threshold. This mask is updated during blending and applied during the initialization of the status map $S$ for the next Gaussian. Any block with $T_{\text{mask}}=1$ is excluded from future $\alpha$ computation.
We found that the termination threshold used in the original 3DGS algorithm ($T<0.0001$) was overly conservative. In our experiments, relaxing this threshold to $T<0.001$ did not lead to any visible degradation in rendering quality. This optimization improves performance while maintaining visual fidelity. The updated $T$ and accumulated color values are written back to the buffer and reused in the blending of subsequent Gaussians.
@@ 4.6. Scheduling under Compatibility Modes
When the hardware detects that the target image exceeds the supported resolution, the pipeline automatically switches to Cmode. In this mode, the image is partitioned into multiple sub-views of manageable size and processed sequentially. To support this, we implement a 2D spatial binning scheme, where Gaussians are grouped not only by depth but also by their screen-space coordinates. Each spatial group corresponds to a specific sub-view, ensuring that the working set remains within the limits of on-chip storage. Based on the analysis in Figure 6, we select a sub-view size of $128\times 128$, which introduces negligible redundancy and hardware overhead while maintaining the original Gaussian distribution.
@@ 5.3. Area and Energy Efficiency
Shared Buffer 0.019 3 $2\times 1\times 6$ KB
SH Buffer 0.116 10 $2\times 3\times 8$ KB
Sorted Buffer 0.029 1 $2\times 1\times 1$ KB
Image Buffer 0.872 37 $1\times 4\times 32$ KB
@@ 5.4. Sensitivity Study
Image Buffer Size. The image buffer is a critical design factor that determines the scalability and edge deployability of the GCC architecture. On resource-constrained platforms such as edge devices, the on-chip buffer capacity directly determines whether full Gaussian-wise + Cross-stage Conditional (CC) rendering can be supported. Based on the analysis in Figure 6, we adopt a $128\times 128$ image buffer combined with compatibility mode to support complex scenes efficiently. Notably, rendering accuracy remains unchanged across different sub-view sizes. This is an expected outcome, as our compatibility mode only alters the processing order of Gaussians. As demonstrated in earlier sections, this configuration achieves strong performance while maintaining reasonable area. To further validate this design choice, we explore the design space from a performance perspective. As shown in Figure 13(a), configurations with 128KB and 512KB buffers yield comparable normalized performance. While a larger buffer reduces runtime latency, the area overhead amortization becomes less favorable, leading to a net decline in area-normalized performance. We select 128KB as the default image buffer size to strike a balance between performance and energy efficiency.
Alpha & Blending Array Size. This component reflects a trade-off between compute parallelism and redundancy control. Larger arrays enable higher throughput by processing more Gaussians in parallel, but they may introduce additional redundancy due to coarser boundary checks. In contrast, smaller arrays provide finer-grained control of Gaussian coverage but sacrifice overall throughput. We explore this trade-off through design space analysis. As shown in Figure 13(b), an $8\times 8$ array provides the best performance when normalized by area. Although slightly larger arrays (e.g., $16\times 16$) can reduce rendering time, the resulting increase in DRAM access latency becomes the dominant performance bottleneck and offsets the gains from additional parallelism.
T-Termination Optimization. Our experiments reveal that the default transparency termination threshold used in 3DGS training ($T<0.0001$) is overly conservative. We find that a stricter threshold of $T<0.001$ introduces no perceptible degradation in rendering accuracy or visual quality across all tested scenes. Although this optimization is not the primary driver of performance, it provides an additional $\sim$10improvement in both latency and energy. This suggests that the original models exhibit representation redundancy, and future work may explore adaptive strategies for determining termination thresholds more efficiently.
@@ 6. Related Work
3DGS System-Level Optimization and Specialized Accelerators. As explicit representations such as 3D Gaussian Splatting (3DGS) gain traction, recent works have explored both GPU-based system-level optimizations and ASIC-level hardware designs. GSCore (Lee et al., 2024) is the first dedicated accelerator for 3DGS inference, offering efficient support for Gaussian sorting, projection, and alpha blending, with energy efficiency significantly outperforming GPU baselines. GauSPU (Wu et al., 2024b) proposes the first 3DGS coprocessor targeting dynamic SLAM scenarios, while GSArch (He et al., 2025) addresses data conflicts during 3DGS training by optimizing parallel Gaussian updates. Shimamura et al. (Shimamura et al., 2025) propose FPGA-based vectorization and spatial parallelism for Gaussian projection, and VR-Pipe (Lee et al., 2025) integrates early alpha termination and block merging into GPU rendering pipelines to approximate the performance of dedicated hardware. While recent works (Wang et al., 2024; Hanson et al., 2025; Feng et al., 2025) have also recognized alpha as a metric for reducing computation, these efforts have focused on pruning Gaussian-tile pairings within conventional pipelines; our approach, in contrast, eliminates this pairing stage and introduces a novel, hardware-friendly, alpha-based search algorithm to efficiently ascertain the effective area of each Gaussian within our Gaussian-wise dataflow. In addition to acceleration efforts for neural volume rendering, recent work such as 3D Gaussian Ray Tracing (Moenne-Loccoz et al., 2024) explores fast ray tracing techniques (Glassner, 1989) for particle-based scenes, highlighting the potential of Gaussian representations beyond splatting-style rasterization. Despite significant progress, most existing approaches still follow traditional tile-wise rendering strategies and stage-decoupled execution models. While these designs offer modularity and reusability, they introduce structural inefficiencies—such as redundant memory transfers, unbalanced compute loads, and poor data reuse—which become more pronounced on bandwidth-constrained or energy-limited platforms. In contrast, our work departs from this paradigm by introducing a Gaussian-wise execution model with cross-stage conditional awareness, enabling holistic pipeline-level optimization for 3DGS inference. This architecture effectively improves throughput and energy efficiency, especially in edge deployment scenarios.

### v3

@@ Abstract.
3D Gaussian Splatting (3DGS) has emerged as a leading neural rendering technique for high-fidelity view synthesis, prompting the development of dedicated 3DGS accelerators for resource-constrained platforms. The conventional decoupled preprocessing-rendering dataflow in existing accelerators has two major limitations: 1) a significant portion of preprocessed Gaussians are not used in rendering, and 2) the same Gaussian gets repeatedly loaded across different tile renderings, resulting in substantial computational and data movement overhead. To address these issues, we propose GCC, a novel accelerator designed for fast and energy-efficient 3DGS inference. GCC introduces a novel dataflow featuring: 1) cross-stage conditional processing, which interleaves preprocessing and rendering to dynamically skip unnecessary Gaussian preprocessing; and 2) Gaussian-wise rendering, ensuring that all rendering operations for a given Gaussian are completed before moving to the next, thereby eliminating duplicated Gaussian loading. We also propose an alpha-based boundary identification method to derive compact and accurate Gaussian regions, thereby reducing rendering costs. We implement our GCC accelerator in 28nm technology. Extensive experiments demonstrate that GCC significantly outperforms the state-of-the-art 3DGS inference accelerator, GSCore, in both performance and energy efficiency.
Accelerators, Gaussian Splatting, 3D Rendering
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 58th IEEE/ACM International Symposium on Microarchitecture; October 18–22, 2025; Seoul, Koreaisbn: 978-X-XXXX-XXXX-X/XX/XXccs: Computer systems organization Architecturesccs: Computing methodologies Rasterizationccs: Hardware Logic circuitsccs: Computing methodologies Computer visionfootnotetext: Co-corresponding authors.
@@ 1. Introduction
First, the two-stage “preprocess-then-render” approach introduces redundant Gaussian movements and computations. During the rendering phase, the early termination strategy in $\alpha$-blending results in significantly fewer Gaussians are used in rendering than are preprocessed. This implies that the 3D Gaussians corresponding to those non-rendered 2D ones can be entirely excluded during preprocessing, thereby reducing both data movements and computations. However, in the standard two-stage dataflow, all 3D Gaussians undergo preprocessing regardless of their actual utilization in subsequent rendering, introducing substantial redundancy. For example, our profiling reveals that approximately 40% of runtime in the state-of-the-art GSCore (Lee et al., 2024) accelerator is dedicated to preprocessing 3D Gaussians, while over 60% of the preprocessed Gaussians remain unused during subsequent rendering.
Second, the tile-wise rendering incurs repeated data accesses for the same Gaussians. At the beginning of rendering, all preprocessed 2D Gaussians establish indexing relationships with image tiles based on their spatial coordinates and bounding sizes. The rendering then processes tiles in scanline order: for each tile, it 1) retrieves all associated Gaussians, 2) sorts them by depth, and 3) computes $\alpha$-blending to determine RGB values for every pixel within the tile. For tiles of fixed size, a Gaussian overlapping multiple tiles is repeatedly loaded for each tile’s rendering. This leads to total Gaussian loads far exceeding the number of unique Gaussians, causing significant memory inefficiency.
@@ 3. GCC Dataflow
[…]
@@ Methods
1:Input: Pixel grid $P$, projected Gaussian center $\mu^{\prime}$, effective condition $E(\cdot)$, visited map $S$
@@
[…]
7:while not $Q.empty()$ do
8:  $p\leftarrow Q.pop()$
9:  for each neighbor $q$ of $p$ do
10:   if $q$ is in bounds and $S[q]$ is false then
11:     $S[q]\leftarrow\text{true}$
12:     if $E(q)$ then
13:      $P_{\text{influence}}\leftarrow P_{\text{influence}}\cup\{q\}$
14:      $Q.push(q)$
15:     end if
16:   end if
17:  end for
18:end while
19:return $P_{\text{influence}}$
@@ 4.4. Alpha Unit
In the 3DGS rendering pipeline, each projected Gaussian must compute a transparency weight $\alpha$ for its coverage on the image plane. As shown in Equation  7, $\alpha$ is modeled as the exponential of a negative quadratic form, derived from the pixel’s relative position and the inverse covariance matrix. To efficiently support this computation in hardware, we implement a lookup-table-based EXP unit using piecewise linear approximation. Given the range of meaningful $\alpha$ values lies in [1/255,1)[1/255,1)[ 1 / 255 , 1 ), the corresponding exponent input is constrained to [5.54,0)[-5.54,0)[ - 5.54 , 0 ). Our LUT covers only this interval; inputs below 5.54-5.54- 5.54 are clamped to $\alpha=0$, and those above 0 are saturated to $\alpha=1$. The LUT contains 16 linear segments, ensuring less than 1% approximation error while significantly reducing area and storage overhead. Unlike GSCore (Lee et al., 2024), which suffers from FP16 overflows in EXP operations due to value mapping, our design avoids such issues through fully fixed-point arithmetic in the EXP unit.
@@ 4.5. Blending Unit
Since 3DGS requires strict back-to-front blending, the pipeline enforces ordering at the block level. If a later Gaussian in the sorted sequence attempts to access a block whose previous Gaussian has not yet completed processing, the pipeline stalls until the prior Gaussian finishes, ensuring correctness of blending results. To avoid unnecessary computation, we introduce a transmittance mask ($T_{\text{mask}}$) that disables blending for blocks whose $T$ value has already fallen below a predefined threshold. This mask is updated during blending and applied during the initialization of the status map $S$ for the next Gaussian. Any block with $T_{\text{mask}}=1$ is excluded from future $\alpha$ computation. The updated $T$ and accumulated color values are written back to the buffer and reused in the blending of subsequent Gaussians.
@@ 4.6. Scheduling under Compatibility Modes
When the hardware detects that the target image exceeds the supported resolution, the pipeline automatically switches to Cmode. In this mode, the image is partitioned into multiple sub-views of manageable size and processed sequentially. To support this, we implement a 2D spatial binning scheme, where Gaussians are grouped not only by depth but also by their screen-space coordinates. Each spatial group corresponds to a specific sub-view, ensuring that the working set remains within the limits of on-chip storage. Based on the analysis in Figure 6, we select a sub-view size of $\times 128$ 128, which introduces negligible redundancy and hardware overhead while maintaining the original Gaussian distribution.
@@ 5.3. Area and Energy Efficiency
Shared Buffer 0.019 3 2×1×62$\times$ 1$\times$ 62 × 1 × 6 KB
SH Buffer 0.116 10 2×3×82$\times$ 3$\times$ 82 × 3 × 8 KB
Sorted Buffer 0.029 1 2×1×12$\times$ 1$\times$ 12 × 1 × 1 KB
Image Buffer 0.872 37 1×4×321$\times$ 4$\times$ 321 × 4 × 32 KB
@@ 5.4. Sensitivity Study
Image Buffer Size. The image buffer is a critical design factor that determines the scalability and edge deployability of the GCC architecture. On resource-constrained platforms such as edge devices, the on-chip buffer capacity directly determines whether full Gaussian-wise + Cross-stage Conditional (CC) rendering can be supported. Based on the analysis in Figure 6, we adopt a $\times 128$ 128 image buffer combined with compatibility mode to support complex scenes efficiently. Notably, rendering accuracy remains unchanged across different sub-view sizes. This is an expected outcome, as our compatibility mode only alters the processing order of Gaussians. As demonstrated in earlier sections, this configuration achieves strong performance while maintaining reasonable area. To further validate this design choice, we explore the design space from a performance perspective. As shown in Figure 13(a), configurations with 128KB and 512KB buffers yield comparable normalized performance. While a larger buffer reduces runtime latency, the area overhead amortization becomes less favorable, leading to a net decline in area-normalized performance. We select 128KB as the default image buffer size to strike a balance between performance and energy efficiency.
Alpha & Blending Array Size. This component reflects a trade-off between compute parallelism and redundancy control. Larger arrays enable higher throughput by processing more Gaussians in parallel, but they may introduce additional redundancy due to coarser boundary checks. In contrast, smaller arrays provide finer-grained control of Gaussian coverage but sacrifice overall throughput. We explore this trade-off through design space analysis. As shown in Figure 13(b), an $\times 8$ 8 array provides the best performance when normalized by area. Although slightly larger arrays (e.g., $\times 16$ 16) can reduce rendering time, the resulting increase in DRAM access latency becomes the dominant performance bottleneck and offsets the gains from additional parallelism.
@@
6. Discussion
To enable efficient 3DGS inference on edge devices for wearable applications, we propose the GCC dataflow and its specialized architecture in this work. However, a noteworthy question arises: Can we satisfy the requirements for 3DGS inference in edge computing scenarios by implementing the GCC dataflow directly on GPUs, thereby obviating the need for custom hardware?
To comprehensively evaluate the effectiveness of GCC dataflow on GPUs, we conducted experiments using both cloud-oriented (NVIDIA RTX 3090) and mobile-oriented (NVIDIA Jetson AGX Xavier) GPU platforms. For both the standard dataflow and GCC dataflow implementations, we use the PyTorch framework and apply equal optimization efforts to ensure fair comparison. Results are shown in Figure 15.
Beyond this result, we identify two noteworthy observations. First, rendering time dominates the overall execution time in GPU-based 3DGS inference, making methods like GCC—which primarily target reducing redundant data movement—offer relatively limited performance improvements. This is primarily due to the abundant on-chip cache of GPUs, making 3DGS inference compute-bound. Conversely, in customized architectures with limited on-chip storage (such as GSCore (Lee et al., 2024) with 272 KB on-chip SRAM), data movement has a more significant impact on inference performance, and as the complexity of the scenario increases (e.g., Drjohnson), the proportion of time spent on data movement becomes even greater. In such cases, our GCC dataflow demonstrates a clear advantage over standard dataflow in improving performance. Second, although GCC reduces rendering computation by identifying more compact Gaussian regions, the rendering time actually increases. The reason is that GCC dataflow is “Gaussian-parallel” in our GPU implementation, where a single thread writes a Gaussian’s contribution to multiple pixels. The “many-to-one” pattern causes data races, forcing costly atomic operations for deterministic blending, which disrupts the GPU’s parallel efficiency. However, our GCC accelerator can fully translate the reduction in rendering computation into savings in inference time through carefully designed compute units and pipeline parallel processing methods.
Note that GCC-based 3DGS inference on the Jetson AGX Xavier delivers just 6$\sim$20 FPS in this experiment, well below the widely recognized 90 FPS target (Kerbl et al., 2023; Huang et al., 2025). By designing dedicated architectures instead of directly implementing on existing GPU platforms, we can fully unleash the performance potential of GCC dataflow, achieving inference performance ($>$150 FPS) comparable to—or even surpassing—high-end GPUs.
Figure 15. Breakdown analysis of per-frame execution time for the standard dataflow and our GCC dataflow on GPUs and accelerators. All results are normalized to that of the standard dataflow in each setting.
@@ 7. Related Work
3DGS System-Level Optimization and Specialized Accelerators. Recent works have accelerated 3DGS through both specialized hardware and system-level optimizations. Dedicated hardware efforts include ASICs targeting inference (Lee et al., 2024), dynamic SLAM (Wu et al., 2024b), and training (He et al., 2025), as well as FPGA-based optimizations for pipeline stages like projection (Shimamura et al., 2025). Concurrently, VR-Pipe (Lee et al., 2025) enhance performance on existing GPUs by integrating early alpha termination into the rendering pipeline.
While some algorithmic optimizations also leverage alpha values to reduce computation (Wang et al., 2024; Hanson et al., 2025; Feng et al., 2025), they operate by pruning Gaussian-tile pairings within the conventional dataflow. Beyond splatting-style rasterization (Glassner, 1989), other works have even explored alternative rendering techniques like ray tracing for Gaussian scenes (Moenne-Loccoz et al., 2024), highlighting the representation’s versatility.
Despite this diversity of approaches, a common thread persists: most prior work adheres to the traditional, decoupled tile-wise rendering paradigm. This model, while modular, introduces structural inefficiencies—such as redundant memory transfers and poor data reuse—which are particularly detrimental on resource-constrained platforms. In contrast, our work fundamentally departs from this paradigm by introducing a holistic, Gaussian-wise execution model with cross-stage conditional awareness. This architecture directly targets the aforementioned inefficiencies, enabling significant gains in throughput and energy efficiency for 3DGS inference.
@@ 8. Conclusion
Acknowledgements.
This work was supported in part by National Key R&D Program of China (No.2022ZD0160304), Beijing Natural Science Foundation (No.4254088), and AMD Heterogeneous Accelerated Compute Clusters (HACC) program.